OpenStudy (anonymous):

What percent of 26 is 4? Write an equation

OpenStudy (anonymous):

4/26 * 100 = X X is the percentage

OpenStudy (pixiedust1):

Here is a little trick I learned\[\frac{ is }{ of }=\frac{ percent }{ 100 }\]

OpenStudy (pixiedust1):

What PERCENT (this means percent is going to be n, it's WHAT. We do not know it.) OF 26! This means 26 goes on the bottom. IS 4! That means 4 goes on the top

OpenStudy (pixiedust1):

Now we have \[\frac{ 4 }{ 26 }=\frac{ n }{ 100 }\]

OpenStudy (pixiedust1):

Now, we do cross multiplying. How would we do that?

OpenStudy (pixiedust1):

Do you know how to cross multiply?